# Anyon interactions in the Chern–Simons–Landau–Ginzburg model of the fractional quantum Hall effect

### v1

@@ Abstract
I Introduction
Topological phases of matter in two spatial dimensions support exotic excitations known as anyons: quasi-particles with fractional statistics interpolating between bosons and fermions. Anyons emerge in diverse condensed matter systems, including the fractional quantum Hall effect (FQHE), where electric charge-magnetic flux attachment generates excitations of fractional charge [1, 2, 3], and in spinor Bose–Einstein condensates where vortex braiding realizes anyonic exchange statistics [4]. A particularly striking proposal is anyon superconductivity [5, 6], in which anyons condense to form a novel superconducting state.
@@ I Introduction
II The Ginzburg–Landau model of anyon superconductivity
The CSLG model is described by the superconducting order parameter $\psi:\mathbb{R}^{2+1}\rightarrow\mathbb{C}$, also known as the Higgs field, and an abelian gauge field $\vec{A}=(A_{0},A_{1},A_{2})\in\mathbb{R}^{2+1}$. Associated to the abelian gauge field is the gauge covariant derivative $D_{\mu}=\partial_{\mu}+iqA_{\mu}$, where $q$ is the gauge charge. The gauge field strength is given by the curvature $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. From the field strength we define the magnetic field $B=F_{12}$ and the electric field $E_{i}=F_{0i}$. We will consider the model defined on Minkowski spacetime $\mathbb{R}^{2+1}$, which is endowed with the Minkowski metric $\eta$ and metric signature (+)(+--).
@@ II The Ginzburg–Landau model of anyon superconductivity
In this letter we are only interested in static vortex anyons, these are defined as minimizers of the static energy with non-trivial topological degree. In the GL model, the electric field is absent when considering statics. However, due to the presence of the CS term (2), we see that the electric field does not vanish $E_{i}=F_{0i}=-\partial_{i}A_{0}\neq 0$ and neither does the kinetic term $D_{0}\psi\overline{D_{0}\psi}=q^{2}A_{0}^{2}|\psi|^{2}\neq 0$. So, the static Lagrangian of the model is
$\mathcal{L}_{\textup{static}}=\,$ $\frac{1}{2}(\partial_{i}A_{0})^{2}+\frac{\kappa}{2}\left(A_{0}B+\epsilon_{ij}A_{i}\partial_{j}A_{0}\right)+\frac{1}{2}q^{2}A_{0}^{2}|\psi|^{2}$
We can simplify this a bit by applying an integration by parts,
$\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,\epsilon_{ij}A_{i}\partial_{j}A_{0}=\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,A_{0}B.$ (4)
Hence, the static Lagrangian can be expressed as
@@
Substituting this into the static energy (II) gives an expression for the static energy that is clearly bounded below by 0,
@@
To summarize, vortices in this Ginzburg–Landau model of anyon superconductivity are minimizers of the positive definite static energy (II), such that the electromagnetic scalar potential $A_{0}$ satisfies the non-linear elliptic Gauss constraint (6).
III Anyon superconductivity
In (2+1)-dimensions, adding a Chern–Simons term to the action endows particles with fractional statistics. This means that when you adiabatically braid one particle around another, the many-body wavefunction $\Psi$ picks up a phase
@@ III Anyon superconductivity
In relation to the FQHE, we can identify the filling fraction $\nu$ from the charge-flux relation (17), that is
$\nu=\frac{|Q_{m}|}{q}=\frac{2\pi\kappa}{q^{2}}.$ (18)
Then, we see that the Noether electric charge in terms of the filling fraction is
$Q_{m}=-qN\nu,$ (19)
which can be fractional as the filling fraction $\nu$ can be fractional. Therefore, the charge carried by each anyon is $q\nu$, which can also be fractional. In addition, the spatial part of the Chern–Simons current is given by
$J^{i}=\kappa\epsilon^{ij}E_{j},$ (20)
from which we can identify the Hall conductance $\sigma_{xy}$ with the CS level, $\sigma_{xy}=\kappa=(q^{2}/2\pi)\nu$.
Importantly, as we have shown, the physical Maxwell charge $Q_{e}$ vanishes for localized anyons, such that the system remains globally electrically neutral. In addition to this, both the magnetic flux $\Phi$ and electric charge $Q_{m}$ are conserved in time. To see this, consider the Maxwell equation
$\frac{\textup{d}\vec{B}}{\textup{d}t}=-(\vec{\nabla}\times\vec{E}).$ (21)
By definition, the electric field is the gradient of the potential, $\vec{E}=-\vec{\nabla}A_{0}$, and there are no additional sources. So the time evolution of the magnetic flux is naturally conserved
$\frac{\textup{d}\vec{\Phi}}{\textup{d}t}=\,$ $-\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,(\vec{\nabla}\times\vec{E})=\int_{\mathbb{R}^{2}}\textup{d}^{2}x\,[\vec{\nabla}\times(\vec{\nabla}A_{0})]=0.$ (22)
Since the electric charge $Q_{m}$ and the flux $\Phi$ are related linearly via the magnetic flux-charge relation (17), it follows that the electric charge $Q_{m}$ is also conserved. This also follows naturally from the definition of a Noether charge.
IV Constrained Newton flow
Static vortex anyons are critical points of the static energy (II), so we must solve the associated Euler-Lagrange field equations of the model and also satisfy the Gauss constraint (6). The Euler-Lagrange field equations are obtained by varying the unreduced static energy functional (II) with respect to the Higgs field $\psi$ and the gauge field $(A_{1},A_{2})$. This gives us the static Ginzburg–Landau equations
@@ IV Constrained Newton flow
Our method for obtaining static vortex anyons is carried out using constrained Newton flow. This is a method developed originally in the context of high energy physics [26, 27] and consists of two main features. The first addresses the Gauss constraint (6) by reformulating the constraint as an unconstrained optimization problem. That is, we want to minimize the functional
@@
V Multi-vortex anyon bound states
The inclusion of the CS term has a profound effect on the binding properties of vortices. Consider the conventional GL model with quartic potential
@@ V Multi-vortex anyon bound states
In the CSLG model, this dichotomy is altered. Each vortex carries not only magnetic flux but also a Noether electric charge tied to the flux by Gauss’ law (6). The resulting electrostatic interaction contributes an additional repulsive channel. Numerical minimization of the static energy (II) with constrained Newton flow reveals that for $\kappa>0$ the interaction behavior of vortices can change qualitatively. For small CS coupling $\kappa$, vortices in the $\lambda<1$ regime still experience overall attraction and tend to coalesce, resembling type I behavior. At intermediate $\kappa$, the long-range Higgs-mediated attraction remains dominant, but the electrostatic repulsion becomes significant at short distances. This frustrates collapse into a single axially symmetric multi-vortex and instead favors the formation of spatially separated bound states with negative binding energy. At larger $\kappa$, the repulsion dominates at all scales, even in the nominal type I regime, driving vortices apart and mimicking type II superconductivity. The resulting binding energies of multi-vortex anyons in the three GL parameter $\lambda$ regimes, for various CS couplings $\kappa$, is shown in Fig.  1.
@@
To explain the observation of the hybrid superconductivity in our single-component model, we consider the asymptotic tails of the vortex anyons and their long-range interactions.
VI.1 Single-vortex anyon tail
Let the vacuum (ground state) have modulus $|\psi|=m>0$, and define the Proca mass $m_{A}=qm$, with $q$ the gauge charge. We work in the Coulomb gauge $\partial_{i}A_{i}=0$ and linearize about
$\psi=m+\phi(r),\quad A_{\mu}=0+a_{\mu},$ (31)
with $|\phi|\ll m$ and the unitary gauge (zero phase) for the Higgs field mode. In general, the Taylor expansion of the potential $V(|\psi|)$ about the ground state $|\psi|=m$ is $V(|\psi|)=V(m)+\frac{1}{2}\phi^{2}V^{\prime\prime}(m)+O(\phi^{3})$. This gives the Higgs field a mass $m_{H}=\sqrt{V^{\prime\prime}(m)}=m\sqrt{\lambda}$. In terms of the gauge field mode $a_{\mu}$, the associated magnetic field is $B=\varepsilon_{ij}\partial_{i}a_{j}$ and the electric field is $E_{i}=-\partial_{i}a_{0}$. Then, the static energy functional, linearized about the vacuum (31), is found to be given by
$E_{\textup{lin}}=\,$ $\frac{1}{2}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\left[\phi\left(-\nabla^{2}+m_{H}^{2}\right)\phi\right]+\frac{1}{2}\int_{\mathbb{R}^{2}}\textup{d}^{2}x\begin{bmatrix}B&a_{0}\end{bmatrix}\begin{bmatrix}\left(-\nabla^{2}+m_{A}^{2}\right)&-\kappa\nabla^{2}\\ \kappa&\left(-\nabla^{2}+m_{A}^{2}\right)\end{bmatrix}\begin{bmatrix}B\\ a_{0}\end{bmatrix}$. (32)
To linear order, the Gauss constraint (6) and the static Ginzburg–Landau equation (23) reduce to
$\left(-\nabla^{2}+m_{H}^{2}\right)\phi=\,$ $0,$ (33a)
$\left(-\nabla^{2}+m_{A}^{2}\right)a_{i}=\,$ $\kappa\epsilon_{ij}\partial_{j}a_{0},$ (33b)
$\left(-\nabla^{2}+m_{A}^{2}\right)a_{0}=\,$ $-\kappa B.$ (33c)
It can be seen that the Higgs-amplitude mode $\phi$ decouples at quadratic order, giving a static Klein-Gordon equation. Let us take the curl of (33b), which yields
$(-\nabla^{2}+m_{A}^{2})B=\kappa\nabla^{2}a_{0},$ (34)
and apply the Laplace operator to the linearized Gauss’ law (33c), giving us
$(\nabla^{2}+m_{A}^{2})\nabla^{2}a_{0}=-\kappa\nabla^{2}B.$ (35)
Then, applying the operator $(\nabla^{2}+m_{A}^{2})$ to (34) and using the relation (35), we obtain a scalar decoupled fourth order Helmholtz like equation for the magnetic field,
$\left[\left(-\nabla^{2}+m_{A}^{2}\right)^{2}+\kappa^{2}\nabla^{2}\right]B=0.$ (36)
This linearized equation takes the same form as the linearized field equation for the order parameter in superfluids with fermionic imbalance [35]. We can factorize this operator into complex-conjugate eigenmodes,
$\Delta_{\kappa}\equiv\,$ $\left[\nabla^{4}-\left(2m_{A}^{2}-\kappa^{2}\right)\nabla^{2}+m_{A}^{4}\right]$
$=\,$ $\left(\nabla^{2}-m_{+}^{2}\right)\left(\nabla^{2}-m_{-}^{2}\right),$ (37)
where the two complex-conjugate gauge eigenmodes are
$m_{\pm}=\alpha\pm i\beta,\quad\alpha=\sqrt{m_{A}^{2}-\frac{\kappa^{2}}{4}},\quad\beta=\frac{\kappa}{2},$ (38)
and they satisfy the following relations,
$m_{+}m_{-}=m_{A}^{2},\quad m_{+}^{2}+m_{-}^{2}=2m_{A}^{2}-\kappa^{2}.$ (39)
We also get the exact same operator acting on the electric field, that is $\Delta_{\kappa}\vec{E}=\vec{0}$. Hence, the complex-conjugate eigenmode pair $m_{\pm}$ are the screening masses of the electric and magnetic fields. To see this, we simply look for asymptotic homogeneous modes with $B(r)\propto e^{-mr}$. Equivalently, we set $\nabla^{2}\mapsto m^{2}$ in the operator $\Delta_{\kappa}$, which yields a characteristic equation whose solutions are exactly $m=m_{\pm}$. Thus, instead of two independent gauge coherence lengths, we have one gauge length plus an oscillation scale set by $\beta$, where the two gauge modes share the same decay rate $\alpha$. The corresponding common gauge coherence (screening) length and oscillation length are
$\lambda_{\textup{gauge}}=\frac{1}{\alpha}=\frac{1}{\sqrt{m_{A}^{2}-\kappa^{2}/4}},\quad\lambda_{\textup{osc}}=\frac{2\pi}{\beta}=\frac{4\pi}{\kappa}.$ (40)
So, the difference between the magnetic field $B$ and the electric field $\vec{E}$ asymptotically is an oscillatory phase, not the decay rate.
This disagrees with the result of Paul–Khare [36], where they claim to find a pair of real penetration depths $\lambda_{\pm}=1/M_{\pm}$, with real static screening masses $M_{\pm}=\sqrt{m_{A}^{2}+\kappa^{2}/4}\,\pm\kappa/2$. The static screening masses they obtain is associated to the operator $(-\nabla^{2}+m_{A}^{2})^{2}-\kappa^{2}\nabla^{2}]$, which has the wrong sign for the linearized static theory. These masses were originally obtained by Pisarski–Rao [37]. However, a subtle distinction is that the masses Pisarski–Rao determined are the propagator pole (dynamical gauge) masses $M_{\pm}$, not the static screening masses $m_{\pm}$. They are related to the static screening masses by a Wick rotation $\kappa\mapsto i\kappa$. The penetration depths are related to the static screening masses, not the propagator pole masses. In the simplest of cases (such as the Ginzburg–Landau model with $\kappa=0$), the propagator pole and static screening masses coincide, however this is not true when the Chern–Simons term (2) is included.
Let us think about this from another perspective. Since the magnetic flux $\Phi$ and the electric charge $Q_{m}$ are related via the relation (17), the electric and magnetic sectors are not independent. That is, a fluctuation in the magnetic field $B$ must be accompanied by a fluctuation in the potential $A_{0}$ and, hence, the electric field $\vec{E}=-\vec{\nabla}A_{0}$. This is why the linearized spectrum does not split into two independent real gauge masses. Instead we get a single decay rate and an oscillation frequency, i.e. a damped oscillatory profile for both the magnetic field and the electric field. So the equality of the two gauge coherence lengths is exactly the linearized reflection of the magnetic flux-charge constraint.
To summarize, we find that the single vortex anyon far tails (asymptotic form) take the form
$\phi(r)=\,$ $c_{H}K_{0}(m_{H}r),$ (41a)
$B(r)=\,$ $c_{B}K_{0}(m_{+}r)+c_{B}^{\ast}K_{0}(m_{-}r),$ (41b)
$a_{0}(r)=\,$ $c_{E}K_{0}(m_{+}r)+c_{E}^{\ast}K_{0}(m_{-}r),$ (41c)
where $c_{H}\in\mathbb{R}$ and $c_{B,E}\in\mathbb{C}$. It is clear that the magnetic and electric field are manifestly real, since $m_{-}=m_{+}^{\ast}$ and $K_{0}$ is analytic.
VI.2 Long-range interactions
Now that we have determined the single-vortex far field form, we proceed to determine the long-range interaction energy following the point-particle method [38, 39, 40]. This is done in two parts, we (i) add linear sources to the linearized energy (32) such that the solutions of the corresponding field equations are exactly the single-vortex far fields, and (ii) compute the interaction energy from the on-shell cross term in the linearization.
Let us introduce sources $S_{H}$, $S_{B}$, and $S_{E}$ that couple linearly to $\phi$, $B$, and $a_{0}$, respectively. This gives us additional terms in the linearized energy,
$E_{\textup{source}}=-\int_{\mathbb{R}^{2}}\textup{d}^{2}x\left[S_{H}\phi+S_{B}B+S_{E}a_{0}\right].$ (42)
The sourced linear equations, associated to $E=E_{\textup{lin}}+E_{\textup{source}}$, are determined to be
$(-\nabla^{2}+m_{H}^{2})\phi=\,$ $S_{H},$ (43a)
$(-\nabla^{2}+m_{A}^{2})B-\kappa\nabla^{2}a_{0}=\,$ $S_{B},$ (43b)
$(-\nabla^{2}+m_{A}^{2})a_{0}-\kappa B=\,$ $S_{E}.$ (43c)
As we did previously, we can eliminate the potential mode $a_{0}$ giving a driven fourth-order equation for the magnetic field $B$,
$\Delta_{\kappa}B=(-\nabla^{2}+m_{A}^{2})S_{B}+\kappa\nabla^{2}\,S_{E}.$ (44)
The homogeneous part of this ($\Delta_{\kappa}B=0$) retains the far field asymptotics.
The Green’s function for the static Klein-Gordon equation in 2D is the Bessel function $K_{0}$, that is, it satisfies
$\left(-\nabla^{2}+\mu^{2}\right)K_{0}(\mu r)=2\pi\delta(r).$ (45)
Therefore, we see that the appropriate source for the Higgs amplitude mode is
$S_{H}(r)=c_{H}(-\nabla^{2}+m_{H}^{2})K_{0}(m_{H}r)=2\pi c_{H}\delta(r).$ (46)
In order to realize exactly the far field form (41b) of the magnetic field with real fields, it is convenient to choose the electric source to be $S_{E}(r)=0$ and then we find the magnetic source to be given by
$S_{B}(r)=2\pi\left[c_{B}\frac{(-\nabla^{2}+m_{-}^{2})}{(-\nabla^{2}+m_{A}^{2})}+c_{B}^{\ast}\frac{(-\nabla^{2}+m_{+}^{2})}{(-\nabla^{2}+m_{A}^{2})}\right]\delta(r).$ (47)
We note that the $(-\nabla^{2}+m_{A}^{2})^{-1}$ operator acting on the Dirac $\delta$-function is just the Yukawa kernel with mass $M$. Operationally, (47) is a linear combination of $\delta$ and $\nabla^{2}\delta$. Substituting the magnetic field source (47) into (44) yields
$\Delta_{\kappa}B=2\pi\left[c_{B}(-\nabla^{2}+m_{-}^{\,2})+c_{B}^{\ast}(-\nabla^{2}+m_{+}^{\,2})\right]\delta(r),$ (48)
whose solution is precisely (41b). We can then determine the electric potential $a_{0}$ from Gauss’ law (6),
$a_{0}(r)=-\kappa\left[\frac{c_{B}}{m_{A}^{2}-m_{+}^{2}}K_{0}(m_{+}r)+\frac{c_{B}^{\ast}}{m_{A}^{2}-m_{-}^{2}}K_{0}(m_{-}r)\right],$ (49)
from which we can identify
$c_{E}=-\kappa\frac{c_{B}}{m_{A}^{2}-m_{+}^{2}}.$ (50)
With the linear sources in-hand, we turn to computing the asymptotic interaction energy of well-separated vortices. Let us consider two identical vortices placed at $\vec{X}_{1}$ and $\vec{X}_{2}$, each with sources (46), (47) translated to its center. Let us define the separation distance $R=|\vec{X}_{1}-\vec{X}_{2}|$ and vortex centres $r_{i}=|\vec{x}-\vec{X}_{i}|$. The interaction energy between well-separated sources comes from the cross-terms in the linearization. That is,
$V_{\textup{int}}(R)=\,$ $-\int_{\mathbb{R}^{2}}\textup{d}^{2}\vec{x}\left[S_{H}^{(1)}\phi^{(2)}+S_{B}^{(1)}B^{(2)}\right]$
$=\,$ $-\int_{\mathbb{R}^{2}}\textup{d}^{2}\vec{x}\left[S_{H}(r_{1})\phi(r_{2})+S_{B}(r_{1})B(r_{2})\right].$ (51)
The Higgs channel gives the usual attractive interaction contribution [41]
$V_{H}(R)=\,$ $-\int_{\mathbb{R}^{2}}\textup{d}^{2}\vec{x}\left[2\pi c_{H}^{2}\delta(r_{1})K_{0}(r_{2})\right]$
$=\,$ $-2\pi c_{H}^{2}K_{0}(m_{H}R),$ (52)
and the gauge channel yields
$V_{B}(R)=\,$ $2\pi\int_{\mathbb{R}^{2}}\textup{d}^{2}\vec{x}\delta(r_{1})\left[c_{B}^{2}K_{0}(m_{+}r_{2})+(c_{B}^{\ast})^{2}K_{0}(m_{-}r_{2})\right]$
$=\,$ $2\pi\left[c_{B}^{2}K_{0}(m_{+}R)+(c_{B}^{\ast})^{2}K_{0}(m_{-}R)\right]$
$=\,$ $4\pi\,\textup{Re}\left[c_{B}^{2}K_{0}(m_{+}R)\right].$ (53)
Let us write the screening mass in the form $m_{+}=\alpha+i\beta=\rho e^{i\theta}$ with $\rho=\sqrt{\alpha^{2}+\beta^{2}}=m_{A}$ and $\theta=\arg(m_{+})$. Then, a power series expansion of the Bessel function $K_{0}$ at large separation $R$ yields
$K_{0}(m_{+}R)\sim\sqrt{\frac{\pi}{2m_{A}R}}\,e^{-\alpha R}\,e^{-i(\beta R+\theta/2)}.$ (54)
Hence, the gauge channel interaction energy becomes
$V_{B}(R)=4\pi|c_{B}|^{2}\sqrt{\frac{\pi}{2m_{A}R}}\,e^{-\alpha R}\cos\left(\beta R-\gamma\right),$ (55)
where $\gamma=\theta/2-2\vartheta$ and $\vartheta=\arg(c_{B})$. So, we see that the long-range oscillation behavior of the gauge modes leads to more interesting long-range interactions.
Putting this all together, we find that the interaction energy of a pair of separated vortex anyons is determined by
$V_{\textup{int}}(R)=\,$ $2\pi|c_{B}|^{2}\sqrt{\frac{2\pi}{m_{A}R}}\,e^{-\alpha R}\cos\left(\beta R-\gamma\right)$
@@ VI.2 Long-range interactions
The standard scalar superconducting order parameter contribution remains monotone attractive, as in the regular GL model. However, the gauge term becomes a damped oscillator with envelope $e^{-\alpha R}/\sqrt{R}$, decay rate $\alpha=\tfrac{1}{2}\sqrt{4m_{A}^{2}-\kappa^{2}}$, and oscillation frequency $\beta=\kappa/2$, alternating between attractive and repulsive behavior. If the gauge term dominates at a shorter range than the Higgs term, it can provide a repulsive force initially ($R<\tfrac{\pi}{\kappa}+\tfrac{2\gamma}{\kappa}$) before switching to an attractive force at longer range ($R>\tfrac{\pi}{\kappa}+\tfrac{2\gamma}{\kappa}$), and repeating this behavior in a decaying oscillatory fashion as the separation distance $R$ increases. In contrast to type 1.5 superconductivity in multicomponent superconductors, where the inter-type behavior arises to due competing length scales with $\xi_{1}<\lambda<\xi_{2}$, this long-range oscillatory behavior can give rise to the shorter-range repulsion and long-range attraction essential for hybrid superconductivity.
VI.3 Ginzburg–Landau limit $\kappa\to 0$
As a consistency check, we want to ensure that we recover the Ginzburg–Landau model in the limit $\kappa\to 0$. First of all, we observe that the decay rate becomes $\lim_{\kappa\rightarrow 0}\alpha=m_{A}=qm$ and the oscillatory behavior vanishes, $\lim_{\kappa\rightarrow 0}\beta=0$. Therefore, the magnetic penetration depth is recovered,
$\lim_{\kappa\rightarrow 0}\lambda_{\textup{gauge}}=\frac{1}{\alpha}=\frac{1}{m_{A}},$ (57)
and the complex-conjugate screening masses tend to the single real-valued Proca mass,
$\lim_{\kappa\rightarrow 0}m_{\pm}=\alpha=m_{A}.$ (58)
Furthermore, we recover the long-range interaction energy of the Ginzburg–Landau model [41, 42],
$\lim_{\kappa\rightarrow 0}V_{\textup{int}}(R)=2\pi\left[c_{A}^{2}K_{0}(m_{A}R)-c_{H}^{2}K_{0}(m_{H}R)\right],$ (59)
where $c_{A}^{2}=2|c_{B}|^{2}$.
VII Conclusion
In this work we have examined vortex physics in the Chern–Simons–Landau–Ginzburg model, highlighting how the addition of a CS term profoundly modifies the landscape of superconductivity in two spatial dimensions. Gauss’ law ties each magnetic flux quantum to a compensating Noether electric charge, rendering vortices intrinsically anyonic. This charge-flux attachment mechanism introduces an electrostatic contribution to the vortex interaction and leads to non-trivial screening behavior. Unlike the Ginzburg–Landau model, where vortices exhibit a simple type I/II dichotomy determined by the GL parameter, the CSLG model yields complex-conjugate gauge screening masses. Consequently, the electric and magnetic fields decay with a common penetration depth but acquire an oscillatory phase shift.
@@ VII Conclusion
The emergence of molecular-like vortex anyon clusters suggests that anyon superconductivity may host a much richer set of collective states than previously appreciated. In addition to extending the classification of superconducting phases, these results highlight deep connections between topological gauge theories, fractional statistics, and non-standard superconductivity. Future directions include a systematic study of vortex lattice phases in this model [45], as well as the role of the CS term in dynamical interactions of vortex anyons [46], and the short-range interactions of vortex anyons [47].
Acknowledgments
I would like to thank E. Babaev, A. Talkachov and M. Speight for useful discussions. The author acknowledges funding from the Olle Engkvists Stiftelse through the grant 226-0103.

### v2

# Anyon Bound States and Hybrid Superconductivity

@@ Abstract
preprint: APS/123-QED
Introduction. Topological phases of matter in two spatial dimensions support exotic excitations known as anyons: quasi-particles with fractional statistics interpolating between bosons and fermions. Anyons emerge in diverse condensed matter systems, including the fractional quantum Hall effect (FQHE), where electric charge-magnetic flux attachment generates excitations of fractional charge [1, 2, 3], and in spinor Bose–Einstein condensates where vortex braiding realizes anyonic exchange statistics [4]. A particularly striking proposal is anyon superconductivity [5, 6], in which anyons condense to form a novel superconducting state.
@@
The Ginzburg–Landau model of anyon superconductivity. The CSLG model is described by the superconducting order parameter $\psi:\mathbb{R}^{2+1}\rightarrow\mathbb{C}$, also known as the Higgs field, and an abelian gauge field $\vec{A}=(A_{0},A_{1},A_{2})\in\mathbb{R}^{2+1}$. Associated to the abelian gauge field is the gauge covariant derivative $D_{\mu}=\partial_{\mu}+iqA_{\mu}$, where $q$ is the gauge charge. The gauge field strength is given by the curvature $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. From the field strength we define the magnetic field $B=F_{12}$ and the electric field $E_{i}=F_{0i}$. We will consider the model defined on Minkowski spacetime $\mathbb{R}^{2+1}$, which is endowed with the Minkowski metric $\eta$ and metric signature (+)(+--).
@@
In this letter we are only interested in static vortex anyons, these are defined as minimizers of the static energy with non-trivial topological degree. In the GL model, the electric field is absent when considering statics. However, due to the presence of the CS term (2), we see that the electric field does not vanish $E_{i}=F_{0i}=-\partial_{i}A_{0}\neq 0$ and neither does the kinetic term $D_{0}\psi\overline{D_{0}\psi}=q^{2}A_{0}^{2}|\psi|^{2}\neq 0$. After an integration by parts, the static Lagrangian can be expressed as
@@
Substituting this into the static energy (Anyon Bound States and Hybrid Superconductivity) gives an expression for the static energy that is clearly bounded below by 0,
@@
To summarize, vortices in this Ginzburg–Landau model of anyon superconductivity are minimizers of the positive definite static energy (Anyon Bound States and Hybrid Superconductivity), such that the electromagnetic scalar potential $A_{0}$ satisfies the non-linear elliptic Gauss constraint (4).
Anyon superconductivity. In (2+1)-dimensions, adding a Chern–Simons term to the action endows particles with fractional statistics. This means that when you adiabatically braid one particle around another, the many-body wavefunction $\Psi$ picks up a phase
@@
Importantly, as we have shown, the physical Maxwell charge $Q_{e}$ vanishes for localized anyons, such that the system remains globally electrically neutral. In addition to this, both the magnetic flux $\Phi$ and electric charge $Q_{m}$ are conserved in time. This follows naturally from the definition of a Noether charge.
Constrained Newton flow. Static vortex anyons are critical points of the static energy (Anyon Bound States and Hybrid Superconductivity), so we must solve the associated Euler-Lagrange field equations of the model and also satisfy the Gauss constraint (4). The Euler-Lagrange field equations are obtained by varying the unreduced static energy functional (Anyon Bound States and Hybrid Superconductivity) with respect to the Higgs field $\psi$ and the gauge field $(A_{1},A_{2})$. This gives us the static Ginzburg–Landau equations
@@
Our method for obtaining static vortex anyons is carried out using constrained Newton flow. This is a method developed originally in the context of high energy physics [26, 27, 28] and consists of two main features. The first addresses the Gauss constraint (4) by reformulating the constraint as an unconstrained optimization problem. That is, we want to minimize the functional
@@
Multi-vortex anyon bound states. The inclusion of the CS term has a profound effect on the binding properties of vortices. Consider the conventional GL model with quartic potential
@@
In the CSLG model, this dichotomy is altered. Each vortex carries not only magnetic flux but also a Noether electric charge tied to the flux by Gauss’ law (4). The resulting electrostatic interaction contributes an additional repulsive channel. Numerical minimization of the static energy (Anyon Bound States and Hybrid Superconductivity) with constrained Newton flow reveals that for $\kappa>0$ the interaction behavior of vortices can change qualitatively. For small CS coupling $\kappa$, vortices in the $\lambda<1$ regime still experience overall attraction and tend to coalesce, resembling type I behavior. At intermediate $\kappa$, the long-range Higgs-mediated attraction remains dominant, but the electrostatic repulsion becomes significant at short distances. This frustrates collapse into a single axially symmetric multi-vortex and instead favors the formation of spatially separated bound states with negative binding energy. At larger $\kappa$, the repulsion dominates at all scales, even in the nominal type I regime, driving vortices apart and mimicking type II superconductivity. The resulting binding energies of multi-vortex anyons in the three GL parameter $\lambda$ regimes, for various CS couplings $\kappa$, is shown in Fig.  1.
@@
Hybrid superconductivity. To explain the observation of the hybrid superconductivity in our single-component model, we consider the asymptotic tails of the vortex anyons and their long-range interactions. That is, we expand about the ground state configuration $\{\psi,A_{\mu}\}=\{m+\phi,0+a_{\mu}\}$ and consider field equations linear in the perturbations $\{\phi,a_{\mu}\}$. We find that the magnetic field and electric field must satisfy the same linearized field equations, $\Delta_{\kappa}B=0$ and $\Delta_{\kappa}E_{i}=0$, where
$\Delta_{\kappa}=\left[\left(-\nabla^{2}+m_{A}^{2}\right)^{2}+\kappa^{2}\nabla^{2}\right].$ (24)
This operator can be factorized into complex-conjugate eigenmodes $m_{\pm}$, which are the static screening masses, given by
$m_{\pm}=\sqrt{m_{A}^{2}-\frac{\kappa^{2}}{4}}\pm i\frac{\kappa}{2}.$ (25)
Therefore, we see that the magnetic and electric fields share a common penetration depth $\lambda_{\textup{gauge}}$ but differ by an oscillation frequency $1/\lambda_{\textup{osc}}$, with
$\lambda_{\textup{gauge}}=\frac{1}{\sqrt{m_{A}^{2}-\kappa^{2}/4}},\quad\lambda_{\textup{osc}}=\frac{4\pi}{\kappa}.$ (26)
That is, asymptotically, the magnetic and electric fields have a damped oscillatory profile and share the same decay rate. The real dynamical gauge masses of the theory are obtained from the static screening masses by a Wick rotation $\kappa\mapsto i\kappa$.
The long-range oscillation behavior of the gauge modes leads to more interesting long-range interactions. We find that the interaction energy of a pair of separated vortex anyons is determined by
$V_{\textup{int}}(R)=\,$ $2\pi|c_{B}|^{2}\sqrt{\frac{2\pi}{m_{A}R}}\,e^{-\sqrt{m_{A}^{2}-\frac{\kappa^{2}}{4}}R}\cos\left(\frac{\kappa}{2}R-\gamma\right)$
@@
The standard scalar superconducting order parameter contribution remains monotone attractive, as in the regular GL model. However, the gauge term becomes a damped oscillator with envelope $e^{-\alpha R}/\sqrt{R}$, decay rate $\alpha=\tfrac{1}{2}\sqrt{4m_{A}^{2}-\kappa^{2}}$, and oscillation frequency $\beta=\kappa/2$, alternating between attractive and repulsive behavior. If the gauge term dominates at a shorter range than the Higgs term, it can provide a repulsive force initially ($R<\tfrac{\pi}{\kappa}+\tfrac{2\gamma}{\kappa}$) before switching to an attractive force at longer range ($R>\tfrac{\pi}{\kappa}+\tfrac{2\gamma}{\kappa}$), and repeating this behavior in a decaying oscillatory fashion as the separation distance $R$ increases. In contrast to type 1.5 superconductivity in multiband superconductors, where the hybrid behavior arises to due competing length scales with $\xi_{1}<\lambda<\xi_{2}$, this long-range oscillatory behavior can give rise to the shorter-range repulsion and long-range attraction essential for hybrid superconductivity.
Conclusion. In this work we have examined vortex physics in the Chern–Simons–Landau–Ginzburg model, highlighting how the addition of a CS term profoundly modifies the landscape of superconductivity in two spatial dimensions. Gauss’ law ties each magnetic flux quantum to a compensating Noether electric charge, rendering vortices intrinsically anyonic. This charge-flux attachment mechanism introduces an electrostatic contribution to the vortex interaction and leads to non-trivial screening behavior. Unlike the Ginzburg–Landau model, where vortices exhibit a simple type I/II dichotomy determined by the GL parameter, the CSLG model yields complex-conjugate gauge screening masses. Consequently, the electric and magnetic fields decay with a common penetration depth but acquire an oscillatory phase shift.
@@
The emergence of molecular-like vortex anyon clusters suggests that anyon superconductivity may host a much richer set of collective states than previously appreciated. Recent works have also explored related anyonic and topological mechanisms in two-dimensional systems: anyonic molecules in fractional quantum Hall liquids [38], doped fractional quantum anomalous Hall insulators exhibiting itinerant anyon superconductivity [39], and chiral pseudospin liquids in moiré heterostructures [40]. These studies underscore the growing interest in anyon-mediated condensation and topological gauge responses, providing complementary microscopic contexts for the effective Chern–Simons mechanism and hybrid vortex behavior examined here. In addition to extending the classification of superconducting phases, these results highlight deep connections between topological gauge theories, fractional statistics, and non-standard superconductivity. Future directions include a systematic study of vortex lattice phases in this model [41], as well as the role of the CS term in dynamical interactions of vortex anyons [42], and the short-range interactions of vortex anyons [43].
Acknowledgments. I would like to thank E. Babaev, A. Talkachov and M. Speight for useful discussions. The author acknowledges funding from the Olle Engkvists Stiftelse through the grant 226-0103.